OpenStudy (anonymous):

Which of the following is the radical expression of a to the four ninths power ?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

someone plz help D;

OpenStudy (zzr0ck3r):

\[a^{\frac{4}{9}}\]?

OpenStudy (zzr0ck3r):

\[a^{\frac{b}{c}}=\sqrt[c]{a^b}\]

OpenStudy (anonymous):

completely lost .. ?@zzr0ck3r

OpenStudy (anonymous):

C would be 9, B would be 4 ?

OpenStudy (anonymous):

and A would stay as it ?? @zzr0ck3r

OpenStudy (zzr0ck3r):

is your question \[\text{"Which of the following is the radical expression of }\large a^{\frac{4}{9}} \]

OpenStudy (anonymous):

yes it is @zzr0ck3r

OpenStudy (anonymous):

the options are.. 4a9 9a4 fourth root of a to the ninth power ninth root of a to the fourth power

OpenStudy (zzr0ck3r):

\[a^{\frac{b}{c}}=\sqrt[c]{a^b}\] just rewrite it like that

OpenStudy (anonymous):

I got D as the answer is the correct ?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thankyou !